# AltFS: Agency-light Feature Selection with Large Language Models in Deep Recommender Systems

### v1

@@ 4 Related Work
Feature Selection for Recommender Systems. Features are a crucial component of machine learning algorithms, and the quality of the feature input directly impacts the model’s predictive performance Liu et al. (2019, 2023); Fan et al. (2020). Therefore, feature selection, an approach aimed at identifying predictive features from the original feature set, has become a critical part of the machine learning pipeline. Traditional feature selection methods can generally be categorized into three types: 1) Filter Methods. These methods define specific criteria for evaluating the importance of features, such as the Chi-squared test Liu and Setiono (1995) and mutual information Battiti (1994). 2) Wrapper Methods. These methods typically employ heuristic algorithms and use black-box models to assess the effectiveness of feature subsets Kohavi and John (1997). 3) Embedded Methods. This type integrates feature selection within the predictive model, evaluating features along with the model optimization process Tibshirani (1996); Friedman (2001). With the advancement of deep learning technologies, numerous related works have emerged in the field of recommender systems, leading to the development of various feature selection methods Li et al. (2024). AutoField Wang et al. (2022) was the first to introduce the use of AutoML for learning feature importance in recommender systems. It constructed an innovative controller network that learns importance weights for different feature fields. AdaFS Lin et al. (2022) expanded on AutoField by recognizing that different features may have varying importance across different samples.
@@
AutoML in Recommender Systems. In recommender systems, challenges such as feature selection, feature crossing, and embedding dimension design face the difficulty of large candidate sets and the need for specialized knowledge to optimize them. AutoML is a technology that effectively lowers the barrier for users by automating and simplifying the design and optimization of machine learning algorithms He et al. (2021). For feature selection, methods like AutoField Wang et al. (2022), AdaFS Lin et al. (2022), and OptFS Lyu et al. (2023) have achieved significant improvements in model performance and efficiency by designing gating weights to automatically select features. In the area of feature crossing, AutoCross Luo et al. (2019) uses beam search in a tree-based space to automatically select high-level cross-features. Regarding embedding dimension search, AutoDim Zhao et al. (2021) employs a two-stage approach: it first conducts end-to-end training of the weights for each embedding dimension in a soft manner, then selects the embedding size with the highest weight for each feature field. Additionally, AutoDis Guo et al. (2021) introduces three core modules—meta-embeddings, automatic discretization, and aggregation—to learn the embedding of features.
5 Conclusion
@@
In this paper, we introduce AltFS, an agency-light feature selection method for deep recommender systems. Specifically, we first propose a prompt iteration technique, which iteratively prompts large language models (LLMs) to obtain an initial ranking of features in the semantic space. Then, we introduce the bridge vector, which connects the semantic information of features with the current recommendation task, further fine-tuning the feature importance ranking. We conduct extensive experiments on three public datasets, and the results demonstrate the superiority of AltFS compared to other state-of-the-art methods.
6 Limitations
AltFS exhibits two potential limitations that are worth addressing in future work. First, although AltFS has optimized computational efficiency, the inference time of LLMs can still become a bottleneck when dealing with datasets containing an extremely large number of features. To mitigate this, future work could explore model quantization or distillation techniques to further enhance efficiency. Second, AltFS relies on the world knowledge encoded in LLMs, which may not fully align with the specific requirements of certain recommendation tasks. Although we have proposed the bridge vector to mitigate this issue, there is still an inevitable gap. To address this, future research could incorporate domain-specific knowledge bases or fine-tune LLMs to better adapt to specific tasks.
@@ Appendix A Appendix
A.1 Prompt
In this section, we will detail the prompt we used in the feature importance extraction stage. The specific prompt we used for Movielens-1M is shown in Figure 6.
Figure 6: Prompt for Movielens-1M.
A.2 Dataset Description
Table 4: Dataset statistics
Dataset Movielens-1M Aliccp Kuairand
Interactions 1,000,209 85,316,519 1,436,609
Users 6,040 238,635 27,285
Items 3,706 467,298 7,551
Interaction Type Rating (1-5) Click Click
Feature Num 9 23 96
@@ A.2 Dataset Description
Figure 7: Hyperparameter analysis on number of LLMs ($K$) and maximum masking ratio ($\beta$) on AUC and Logloss for Aliccp dataset.
A.3 Baseline Introduction
In this section, we will provide a detailed introduction to the baselines.
A.3.1 Shallow Feature Selection Methods.
  • Lasso Tibshirani (1996). Lasso is the Least Absolute Shrinkage and Selection Operator, which effectively combines variable selection and regularization to enhance model performance.
  • GBDT Friedman (2001). Gradient Boosted Decision Trees (GBDT) is a highly effective machine learning algorithm that leverages the strengths of decision trees in combination with the boosting technique. It constructs an ensemble of decision trees, with each tree iteratively correcting the errors of its predecessors.
  • Random Forest Breiman (2001). Random Forest determines feature importance by measuring the extent to which each feature reduces impurity within a decision tree, typically using Gini impurity or entropy.
  • XGBoost Chen and Guestrin (2016). XGBoost is a highly efficient and scalable machine learning algorithm built on the principles of gradient boosting. XGBoost determines feature importance by assessing the impact of each feature on the model’s final predictive performance.
A.3.2 Gate-based Feature Selection Methods.
  • AutoField Wang et al. (2022). AutoField is the first work in the recommender systems domain to focus on feature selection within DRS (Deep Recommender Systems). It introduces an advanced controller network that trains two parameters for each feature field, representing the likelihood of selecting that feature.
  • AdaFS Lin et al. (2022). AdaFS refines feature selection further along the lines of AutoField. Recognizing that feature importance varies significantly across different samples, AdaFS focuses on generating dynamic gating weights for each sample, enabling the model to adaptively select features.
  • OptFS Lyu et al. (2023). OptFS differs from previous work by focusing on feature value selection. Through the learning of a gating network, OptFS can significantly reduce the number of feature values while simultaneously enhancing the model’s predictive performance.
  • LPFS Guo et al. (2022). LPFS critiques previous approaches that determine feature importance based on the magnitude of gating values. It introduces a smoothed-$l^{0}$ function that effectively selects features in a single stage.
A.3.3 Sensitivity-based Feature Selection Methods.
  • Permutation Fisher et al. (2019). This method assesses the importance of a feature by disrupting the values within a specific feature field and observing whether this permutation significantly impacts the model’s predictive performance.
  • SFS Wang et al. (2023). SFS assigns feature importance weights by examining the gradient of the gating vector on a small subset of data.
  • SHARK Zhang et al. (2023). SHARK uses the first-order component of the Taylor expansion as a measure of feature importance. It improves model efficiency and performance by removing less important features from the embedding table.
A.4 More details on implementation
In this section, we will introduce more details on implementation to ease reproduction.
A.4.1 Hyperparameter settings
For AltFS, we selected three LLMs: GPT-4, GPT-4o, and GPT-3.5. In the hyperparameter analysis, we also explored the performance when using four LLMs, with the additional LLM being Gemini-1.5-Pro. Our experiments were conducted on a single A100 GPU. The hyperparameter settings for other baselines are detailed in Table 5.
Table 5: Hyperparameters for baselines.
@@ A.4.1 Hyperparameter settings
A.4.2 Guidelines for reproduction
In this section, we primarily explain how to reproduce the results. The information is also included in the readme file in our code repository.
Feature Importance Extraction stage. You can run python llm_predict/predict.py –model=gpt-4-turbo –dataset=movielens to generate feature importance sequences by LLMs. The generating sequences will be saved automatically.
Feature Importance Refinement stage. python fs_run.py –dataset=movielens-1m –fs=altfs –retrain=false This process will run the Feature Importance Refinement stage. The terminal will output the will-trained bridge vector when the process ends.
Retraining stage. Running script by python llm_predict/llm_selection.py to get the final feature rank file. Then running python fs_run.py –dataset=movielens-1m –model=dcn –train_or_search=false –rank_path=altfs_movielens-1m –k=k –checkpoint_folder=checkpoints_for_retrain/ to retrain the recommendation model.
A.5 More results on Hyperparameter analysis
In this section, we will give the hyperparameter analysis results on both AUC and Logloss for Aliccp dataset. Since a higher AUC is better and a lower Logloss is preferable, we can observe from Figure 7 that parameter with higher AUC values correspond to lower Logloss values, indicating an inverse relationship between the two metrics. This demonstrates that our analysis of AUC in the hyperparameter analysis is equally applicable to Logloss.
Number of LLMs: Experimenting with 1 to 4 LLMs, the results reveal that a single strong LLM (e.g., GPT-4) is beneficial, outperforming the no-selection baseline. Optimal performance is achieved with 3 LLMs, as the bridge vector effectively integrates feature importance sequences from different LLMs, balancing semantic and task-specific aspects. Using more than 3 LLMs stabilizes performance, with 3 being the most cost-effective choice.
Maximum Masking Ratio: Varying the maximum masking ratio from 0.0 to 1.0 shows an initial increase in AUC followed by a decrease and an initial decrease in Logloss followed by an increase. A lower ratio (near 0) averages LLM-generated sequences, while a higher ratio (near 1) reduces the differences among features, weakening selection effectiveness. The performance remains robust between ratios of 0.2 and 0.5, indicating the stability and generalization ability of AltFS.
A.6 Case Study
Figure 8: (a) Cramér’s V matrix and (b) AUC performance on Movielens-1M.
In this section, we present a specific example to illustrate the effectiveness of LLM reasoning in the feature selection process. Figure 8(a) shows the Cramér’s V index of various features in the MovieLens-1M dataset, representing the correlation between different features. The darker the green, the stronger the correlation between the two features. We observe that “title” and “movie_id” are highly correlated, as indicated by the red box in the figure. Additionally, Figure 8(b) displays the AUC performance on the MovieLens-1M dataset for the best-performing baseline LPFS, LPFS-t (which removes the title feature based on LPFS), and AltFS. We find that removing the title feature results in a significant improvement for LPFS, indicating that the strong correlation between “title” and “movie_id” indeed hinders model training. Finally, we examine the feature importance ranking from LPFS: [“movie_id”, “user_id”, “title”, “genres”, “zip”, “gender”, “age”, “occupation”, “timestamp”]. LPFS assigns high importance to both “movie_id” and “title”, likely because the model optimization process treats these two features as roughly equally important, overlooking the issue of information redundancy between them. However, AltFS, by leveraging LLM reasoning, effectively identifies this issue. Below is an analysis extracted from the LLM response during prompt iteration regarding the title feature: “While the movie title might have some influence, it’s likely highly correlated with movie_id and thus doesn’t add significant independent information.” Consequently, a more accurate and effective feature ranking is obtained: [“user_id”, “movie_id”, “genres”, “age”, “gender”, “timestamp”, “occupation”, “zip”, “title”]. This demonstrates the effectiveness of AltFS in handling feature importance extraction.

### v2

@@ 4. Related Work
Feature selection, an approach aimed at identifying predictive features from the original feature set, has become a critical part of the machine learning pipeline (Liu et al., 2019, 2023; Fan et al., 2020; Li et al., 2025; Jia et al., 2024a) and data mining framework (Jia et al., 2025; Zhang et al., 2025; Jia et al., 2024e, b, 2023). Traditional feature selection methods can generally be categorized into three types: 1) Filter Methods. These methods define specific criteria for evaluating the importance of features, such as the Chi-squared test (Liu and Setiono, 1995) and mutual information (Battiti, 1994). 2) Wrapper Methods. These methods typically employ heuristic algorithms and use black-box models to assess the effectiveness of feature subsets (Kohavi and John, 1997). 3) Embedded Methods. This type integrates feature selection within the predictive model, evaluating features along with the model optimization process (Tibshirani, 1996; Friedman, 2001). With the advancement of deep learning technologies, numerous related works have emerged in the field of recommender systems, leading to the development of various feature selection methods (Li et al., 2024). AutoField (Wang et al., 2022) was the first to introduce the use of AutoML for learning feature importance in recommender systems. It constructed an innovative controller network that learns importance weights for different feature fields. Recently, there have been some attempts (Jeong et al., 2024) to use LLMs for feature selection, but the recommendation system context has not been considered.
4.2. AutoML in Recommender Systems.
@@
In recommender systems (Jia et al., 2018, 2024d), challenges such as feature selection, feature crossing, and embedding dimension design face the difficulty of large candidate sets and the need for specialized knowledge to optimize them. AutoML is a technology that effectively lowers the barrier for users by automating and simplifying the design and optimization of machine learning algorithms (He et al., 2021). For feature selection, methods like AutoField (Wang et al., 2022), AdaFS (Lin et al., 2022), and OptFS (Lyu et al., 2023) have achieved significant improvements in model performance and efficiency by designing gating weights to automatically select features. In the area of feature crossing, AutoCross (Luo et al., 2019) uses beam search in a tree-based space to automatically select high-level cross-features.
5. Conclusion
@@
In this paper, we introduce SELF, an agency-light feature selection method for DRS. Specifically, we first propose a prompt iteration technique, which iteratively prompts LLMs to obtain an initial ranking of features in the semantic space. Then, we introduce the bridge vector, which connects the semantic information of features with the current recommendation task, further fine-tuning the feature importance ranking. We conduct extensive experiments on three public datasets, and the results demonstrate the superiority of SELF compared to other state-of-the-art methods. We also release our code online for ease of reproduction.
Acknowledgements.
This research was partially supported by Hong Kong Research Grants Council’s Research Impact Fund (No.R1015-23), Collaborative Research Fund (No.C1043-24GF), General Research Fund (No.11218325), Institute of Digital Medicine of City University of Hong Kong (No.9229503), and Huawei (Huawei Innovation Research Program).
@@ Appendix A Appendix
A.1. Dataset Description
@@
A.2. More details on implementation
In this section, we will introduce more details on implementation to ease reproduction. For SELF, we selected three LLMs: GPT-4, GPT-4o, and GPT-3.5. In the hyperparameter analysis, we also explored the performance when using four LLMs, with the additional LLM being Gemini-1.5-Pro. Our experiments were conducted on a single A100 GPU. The hyperparameter settings for other baselines are detailed in Table 5.
Table 5. Hyperparameters for baselines.
@@ A.2. More details on implementation
GenAI Usage Disclosure
Generative AI tools were used only for spelling and grammar correction purposes. No content was generated or modified beyond these basic editing tasks.